# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v2

@@ 5.2. Datasets
The LQA datasets are a specialised resource designed to facilitate research in the domain of legal NLP. they consist of a collection of legal questions and corresponding answers, drawn for various legal documents and case law. Most questions in the LQA datasets fall into two main categories: knowledge-driven questions (KD-questions) and case-analysis questions (CA-questions) (Zhong et al., 2020c). KD-questions are centred around the understanding of specific legal concepts, whereas CA-questions involve the analysis of actual legal cases. Both types demand advanced reasoning skills and a deep comprehension of the text, making LQA a particularly challenging task in the field of NLP.
Zhong et al. (2020c) introduce JEC-QA, a dataset with 26,365 multiple-choice questions from the National Judicial Examination of China and related websites. Each question offers four possible answers and is labelled with the type of reasoning required, such as word matching, concept understanding, numerical analysis, multi-paragraph reading, and multi-hop reasoning. This dataset poses challenges for QA models, highlighting the gap between machine performance and human expertise in legal reasoning.
Sovrano et al. (2021) present a dataset designed to evaluate automated QA systems within the domain of Private International Law. It includes 17 carefully selected questions based on key EU regulations—Rome I, Rome II, and Brussels I bis—with answers derived directly from these regulations. The questions are classified based on their specificity, allowing for nuanced analysis of context-dependency in legal reasoning. This dataset aids in assessing the performance of QA systems intended for legal professionals navigating complex cross-border legal issues.
Büttner and Habernal (2024) introduce GerLayQA, a dataset supporting LQA for laypersons in Germany, focusing on the civil-law system. It contains 21,538 real-world questions posed by laypersons, paired with expert answers from lawyers grounded in specific paragraphs of German law books. The dataset are constructed through filtering and quality assurance to ensure accuracy and relevance, making it a valuable resource for developing LQA systems that can interpret and apply German law to everyday legal inquiries.
6. Legal Judgement Prediction
LJP is an important task within legal NLP, especially prevalent in civil law systems where judgements are determined based on case facts and statutory articles (Zhong et al., 2020b). This task involves predicting legal outcomes from the descriptions of cases and relevant legal statutes, and is essential in countries, such as France, Germany, Japan, and China (Zhong et al., 2020b). LJP has attracted interest from AI researchers and legal professionals due to its potential to assist judges, lawyers, and legal scholars in predicting case outcomes based on historical data (Chen et al., 2024a).
Despite its promise, LJP is a highly demanding and challenging task. It requires careful handling of natural biases in historical legal data, which can create feedback loops that amplify discrimination (Krasadakis et al., 2024). Therefore, ensuring the impartiality of predicted rulings is crucial (Krasadakis et al., 2024). Currently, LJP is primarily performed by legal experts who undergo extensive specialised training to manage the complex steps involved, such as identifying relevant law articles, defining charge ranges, and deciding penalty terms (Cui et al., 2023). Nevertheless, LJP provides substantial benefits, streamlining legal decision-making processes for both practitioners and ordinary citizens (Cui et al., 2023).
6.1. Task
@@
Zhong et al. (2018) introduce TopJudge, a topological MTL framework that models dependencies among subtasks in LJP, such as law article prediction, charge prediction, and penalty terms. Using a Directed Acyclic Graph, TopJudge processes subtasks in a topological order reflecting real-world legal decision-making. Evaluated on large-scale Chinese criminal case datasets, it outperformed previous models in predicting legal outcomes.
@@
Ye et al. (2018) address the problem of Court View Generation from fact descriptions in criminal cases to enhance the interpretability of charge prediction systems and aid in automatic legal document generation. They formulated this as a text-to-text Natural Language Generation (NLG) problem, using a label-conditioned Seq2Seq model with attention to decode court views based on encoded charge labels. This work contributes to automatic legal document generation by providing justifications for charge decisions.
@@
Yang et al. (2019) propose a Multi-Perspective Bi-Feedback Network (MPBFN) with a Word Collocation Attention mechanism to improve LJP. MPBFN addresses the challenges of multiple subtasks and their dependencies by using a bi-feedback mechanism for forward prediction and backward verification among subtasks. The Word Collocation Attention integrates word collocation features and numerical semantics to better predict penalties. Evaluated on the CAIL datasets (Xiao et al., 2018), their model outperformed baselines in predicting law articles, charges, and penalty terms.
@@
Chalkidis et al. (2019a) introduce an English LJP dataset containing approximately 11.5k cases from the European Court of Human Rights (ECHR). They evaluated various neural models on this dataset, including a hierarchical version of BERT (HIER-BERT) to handle long legal documents. Their models outperformed previous feature-based approaches in tasks, such as violation classification and case importance prediction. They also explored potential biases in legal predictive models using data anonymization.
@@
Medvedeva et al. (2020) investigate using NLP tools to predict judicial decisions of the ECHR based on court proceeding texts. They employed Support Vector Machine (SVM) linear classifier to predict violations of articles, achieving an average accuracy of 75%. However, when predicting future decisions based on past cases, accuracy decreased to 58–68%. They also found that predicting outcomes based solely on judges’ surnames could achieve accuracy of 65%, highlighting potential biases.
@@
Zhong et al. (2020a) introduce QAjudge, a reinforcement learning-based model designed to provide interpretable legal judgements by visualising the prediction process. QAjudge uses a Question Net to iteratively select relevant yes-no questions about case facts, an Answer Net to provide answers, and a Predict Net to generate the final judgement. The model aims to minimise the number of questions asked. Evaluated on real-world datasets, QAjudge demonstrated potential in providing reliable and transparent legal judgements.
@@
Xu et al. (2020) propose the Law Article Distillation based Attention Network (LADAN), an end-to-end model addressing the issue of confusing charges in LJP by distinguishing similar law articles. The model uses a novel graph NN to learn differences between confusing law articles and an attention mechanism to extract discriminative features from fact descriptions. Experiments on real-world datasets showed that LADAN improved performance over previous methods in law article prediction, charge prediction, and penalty term prediction.
@@
Ma et al. (2021) introduce MSJudge, a MTL framework designed to predict legal judgements by leveraging multi-stage judicial data, including pre-trial claims and court debates. MSJudge consists of components to encode multi-stage context, model interactions among claims, facts, and debates, and predict judgements. Evaluated on a large civil trial dataset, MSJudge outperformed state-of-the-art baselines in judgement quality.
@@
Feng et al. (2022) address limitations of state-of-the-art LJP models by proposing an event-based prediction model with constraints to improve performance. The model extracts fine-grained key events from case facts and predicts judgements based on these events rather than the entire fact statement. They manually annotated a legal event dataset and introduced output constraints to guide learning. Their method leverages event information and cross-task consistency constraints.
@@
Tong et al. (2024) introduce GJudge, a graph boosting framework incorporating constraints to address shortcomings of traditional LJP methods. GJudge features a multi-perspective interactive encoder and a Multi-Graph Attention Network (MGAT) consistency expert module. The encoder merges fact descriptions with label similarity connections, while the expert module distinguishes similar labels and maintains task consistency. Testing on datasets showed that GJudge outperformed other models, including the state-of-the-art RLJP (Wu et al., 2022) model, with higher F1 scores.
@@
Previous works mainly focus on creating accurate representations of a case’s fact description to enhance judgement prediction performance. However, these methods often overlook the practical judicial process, where human judges compare similar law articles or potential charges before making a decision. To address this gap, Zhang et al. (2023a) propose CL4LJP, a supervised contrastive learning framework to improve LJP by capturing fine-grained differences between similar law articles and charges. The framework includes contrastive learning tasks at the article, charge, and label levels, enhancing the model’s ability to model relationships between fact descriptions and labels.
@@
Liu et al. (2023) propose ML-LJP, a Multi-Law aware LJP method that expands law article prediction into a multi-label classification task incorporating both charge-related and term-related articles. The approach uses label-specific representations and contrastive learning to distinguish similar definitions. A Graph Attention Network (GAT) is employed to learn interactions among multiple law articles for prison term prediction. Experiments showed that ML-LJP outperformed state-of-the-art models, particularly in prison term prediction.
6.2. Datasets
The LJP datasets are specialised resources designed to advance research in predicting judicial outcomes within the domain of legal NLP. These datasets are categorised into four main types: court view generation datasets, law articles, charge prediction, and prison term prediction. Court view generation datasets involve court opinions and summaries. Law articles datasets involve the prediction of legal outcomes based on specific statutes or regulations. Charge prediction datasets are concerned with predicting the charges that should be brought against a defendant based on the case details, while prison term prediction datasets aim to estimate the likely sentence duration given the nature of the crime and the legal context. Each type of dataset presents unique challenges, demanding not only text comprehension but also the ability to apply complex legal reasoning.
Court View Gen (Ye et al., 2018) is an innovative dataset containing 171,981 Chinese legal cases, each involving a single defendant and a corresponding charge, covering a total of 51 charge categories. This dataset is specifically curated to support the generation of court opinions based on charge labels. The data was sourced from publicly available legal documents within the CJO [^3] http://wenshu.court.gov.cn repository.
Semo et al. (2022) introduce the first LJP dataset focused on class action cases in the US. The dataset targets predicting outcomes based on plaintiffs’ complaints rather than court-written fact summaries, involving a rule-based extraction system to identify relevant text spans from complaints.
@@ 7. Legal Text Classification
LTC is an important task within the domain of NLP that involves categorising legal documents based on their content, a foundational aspect of building intelligent legal systems. With the exponential growth of legal documents, it has become increasingly challenging for legal professionals to locate relevant rulings in similar cases for argumentation. LTC addresses this challenge by automatically associating legal texts with predefined categories, such as criminal, civil, or administrative cases, thereby simplifying legal research and decision-making processes.
@@
In the legal field, this process is often referred to as predictive coding, where ML algorithms are trained through supervised learning to classify documents into specific categories. The broader task of text classification in NLP involves assigning one or multiple categories to a document from a set of predefined options, and it can take various forms, including binary classification, multi-class classification, and multi-label classification. Legal document classification often falls under large multi-label text classification, where the label space can consist of thousands of potential categories, adding complexity to the task (Shaheen et al., 2020).
7.1. Task
DL typically requires extensive data to yield effective results, but MTL offers a potential solution to the data scarcity problem. Elnaggar et al. (2018) leverage transfer learning and MTL to perform tasks, such as translation and multi-label classification within legal document corpora. They utilise the MultiModel algorithm (Kaiser et al., 2017), which uses a fully convolutional sequence-to-sequence architecture integrating different modality nets. Their model processes legal texts by mapping them into a shared embedding space, allowing task switching and supporting generalisation across legal tasks, thereby addressing data scarcity in the legal field.
Wei et al. (2018) investigate the application of CNNs for text classification in legal document review, comparing their performance with SVMs on four datasets from real legal cases. They found that CNNs perform better with larger training datasets and offer a more stable growth trend compared to SVMs. The study highlighted challenges such as sequence length limitations and the need for explainability in predictive analysis, suggesting improvements, such as integrating annotated sentences with full text to enhance sentence relevance identification.
Lee and Lee (2019) focus on LTC in the Korean language and compare three different DL approaches: CNN with ASCII encoding, CNN with Word2Vec embeddings, and RNN with Word2Vec embeddings. The classification models are used to classify case data into civil, criminal, and administrative. Using a dataset of nearly 60,000 past case documents, the study finds that the RNN model with Word2Vec embedding achieves the highest classification accuracy.
Bambroo and Awasthi (2021) introduce an architecture that integrates long attention mechanisms with a distilled BERT model pre-trained on legal domain-specific corpora. Their model employs a combination of local windowed attention and task-motivated global attention to handle inputs up to eight times longer than standard BERT models. The architecture, based on DistilBERT (Sanh et al., 2020) and incorporating LongformerSelf-Attention, is optimised for legal document classification, outperforming fine-tuned BERT and other transformer-based models in both speed and performance.
Song et al. (2022) present a deep learning-based system built on top of RoBERTa (Liu et al., 2019) for multi-label legal document classification. They enhance the model with domain-specific pre-training, a label-attention mechanism, and MTL to improve classification accuracy, particularly for low-frequency classes. The label-attention mechanism uses label embeddings to bridge the semantic gap between samples and class labels, addressing class imbalance issues.
Wang et al. (2022) introduce a Document-to-Graph Classifier to classify legal documents based on facts and reasons rather than topics. They extract key entities and represented legal documents using four distinct relation graphs capturing different aspects of entity relationships. A GATs (Veličković et al., 2018) is used to learn document representations from the combined graph, improving classification by focusing on factual content.
Mamooler et al. (2022) propose an active learning pipeline for fine-tuning PLMs for LTC. They address challenges of specialised vocabulary and high annotation costs. Their method involves continued pre-training of RoBERTa on legal texts, knowledge distillation using a pre-trained sentence transformer, and an initial sampling strategy by clustering unlabelled data. This approach reduces the number of labelling actions required for LTC tasks.
Grabmair et al. (2015) introduce LUIMA, a system designed for conceptual legal document retrieval, focusing on vaccine injury decisions. LUIMA employs a multi-level classification pipeline: rule-based sub-sentence annotations tag legal concepts, such as terms and mentions, while machine learning classifiers categorise sentences into argumentative roles, such as legal rules or evidence-based findings. These annotations feed into a sentence-level indexing process using Apache Lucene, enabling semantic querying beyond traditional keyword search. A learning-to-rank module refines the retrieved results by leveraging features, such as sentence match counts and similarity scores.
Galassi et al. (2024) explore the binary classification task of detecting potentially unfair clauses in online Terms of Service (ToS) at the sentence level. Sentences are labelled as either unfair/potentially unfair or fair. Starting from an English-trained model, the authors compare four cross-lingual approaches: (1) training separate models on each language, (2) projecting annotations from English to another language, (3) translating English training documents and using their original annotations, and (4) translating query documents into English during prediction. The evaluation shows scenario (4) achieves similar or better results than scenario (1) when translation quality is high. If the translation quality is low, scenario (1) can produce better results. Scenarios (2) and (3) perform slightly below scenario (1) but remain viable alternatives.
Rhetorical role labelling is the task of classifying sentences in legal documents based on their functional roles, such as fact, argument, or ruling, to structure and analyse judicial decisions. The Artificial Intelligence for Legal Assistance (AILA) shared task series focuses on advancing legal NLP by introducing datasets and challenges that address core legal text processing tasks. The latest edition, AILA 2021 (Parikh et al., 2022), featured a rhetorical role labelling task that required classifying sentences into seven predefined roles. Building on these challenges, DeepRhole (Bhattacharya et al., 2023) introduces a transformer-based framework that fine-tunes domain-adapted models such as LegalBERT to capture the semantics of legal texts. The system applies different word embedding strategies, including Law2Vec and embeddings from Google News, to model legal language variability. An inter-annotator study examines the subjectivity in rhetorical role assignment, further assessing the model’s performance in different judicial contexts.
@@ 7.2. Datasets
Chalkidis et al. (2019b) release EURLEX57K, a dataset containing 57,000 EU legislative documents from the EUR-LEX portal444lex.europa.eu/, annotated with EuroVoc [^5] http://eurovoc.europa.eu/ concepts. This dataset facilitates research in LTC, including extreme multi-label text classification, few-shot, and zero-shot learning, with documents tagged with an expansive set of descriptors.
Tuggener et al. (2020) introduce LEDGAR, a multi-label corpus of legal provisions from contracts scraped from the U.S. Securities and Exchange Commission’s website. The dataset includes over 846,000 provisions across 60,540 contracts, with an extensive label set suitable for text classification and legal studies, aiding in developing legal NLP models.
@@
Papaloukas et al. (2021) introduce the Greek Legal Code dataset, categorising approximately 47,000 Greek legislative documents into a detailed multi-level classification system. The dataset is structured into volumes, chapters, and subjects, each containing diverse legal documents from Greek legislation history, supporting LTC in the Greek legal domain.
@@
Song et al. (2022) introduce POSTURE50K, a legal dataset containing 50,000 U.S. legal opinions annotated with Legal Procedural Postures ranging from common to rare motions. The dataset includes an innovative split strategy to support supervised and zero-shot learning evaluations, ensuring infrequent categories are adequately represented, enhancing model generalizability and testing accuracy.
@@
Graham et al. (2023) develop a domain-specific dataset for LTC focusing on deontic modalities in contract sentences. They manually annotated contract sentences to train models for identifying deontic sentences, such as permissions, obligations, and prohibitions. The corpus, derived from the Contract Understanding Atticus Dataset (CUAD) (Hendrycks et al., 2021), provides a resource for studying functional categories for legal analysis.
@@
Galassi et al. (2024) extend the Claudette [^6] https://claudette.eui.eu corpus from 25 to 50 ToS per language (English, German, Italian, Polish). Each sentence is labelled according to nine categories of potential unfairness, with annotations indicating the degree of unfairness. The dataset was carefully compiled based on language availability, structural similarity, and version correspondence. Cross-lingual analysis revealed notable differences between languages—particularly in German—including variations in length, structure, missing clauses, and legal terminology, reflecting manual drafting adjustments rather than simple automated translations.

### v3

@@ 6.1. Datasets
The LJP datasets are specialised resources designed to advance research in predicting judicial outcomes within the domain of legal NLP. These datasets are categorised into four main types: court view generation, law articles, charge prediction and prison term prediction. Court view datasets contain judicial opinions and summaries. Law articles datasets focus on predicting outcomes based on specific statutes or regulations. Charge prediction datasets aim to identify the charges appropriate to the case details. Prison term datasets estimate likely sentence durations given the crime and legal context. Each category presents unique challenges, requiring not only text comprehension but also the ability to apply complex legal reasoning.
The Court View Gen (Ye et al., 2018) dataset is an innovative resource containing 171,981 Chinese legal cases, each involving a single defendant and a corresponding charge, covering a total of 51 charge categories. It is specifically curated to support the generation of court opinions based on charge labels. All cases were collected from the publicly available China Judgements Online repository.
Semo et al. (2022) introduce the first LJP dataset centred on US class-action lawsuits. Unlike prior work that relies on court-written summaries of facts, this dataset targets outcome prediction directly from plaintiffs’ complaints. A rule-based extraction system identifies the relevant text spans within each complaint.
6.2. Approaches
@@
Zhong et al. (2018) introduce TopJudge, a topological MTL framework that models dependencies among subtasks in LJP, such as law article prediction, charge prediction and penalty terms. Using a directed acyclic graph, TopJudge processes subtasks in a topological order reflecting real-world legal decision-making. Evaluated on large-scale Chinese criminal case datasets, it outperformed previous models in predicting legal outcomes.
@@
Ye et al. (2018) tackle court view generation from criminal case fact descriptions to improve the interpretability of charge prediction systems and support automatic legal document drafting. They frame the task as a text-to-text natural language generation problem, employing a label-conditioned sequence-to-sequence model with attention to generate court views from encoded charge labels. Their approach advances automatic legal document generation by providing explicit justifications for charge decisions.
@@
Yang et al. (2019) propose a Multi-Perspective Bi-Feedback Network (MPBFN) with a Word Collocation Attention mechanism to improve LJP. The MPBFN addresses the challenges of multiple subtasks and their dependencies by using a bi-feedback mechanism for forward prediction and backward verification among subtasks. The Word Collocation Attention integrates word collocation features and numerical semantics to better predict penalties. Evaluated on the CAIL datasets (Xiao et al., 2018), their model outperformed baselines in predicting law articles, charges, and penalty terms.
@@
Chalkidis et al. (2019a) introduce an English LJP dataset containing approximately 11,500 cases from the European Court of Human Rights (ECHR). They evaluated various neural models on this dataset, including a hierarchical version of BERT (HIER-BERT) to handle long legal documents. Their models outperformed previous feature-based approaches in tasks, such as violation classification and case importance prediction. They also explored potential biases in legal predictive models using data anonymization.
@@
Medvedeva et al. (2020) use a linear Support Vector Machine (SVM) to predict whether the ECHR will find a violation of any of nine Convention articles. Trained on the textual proceedings, the model achieves an average 75% accuracy, yet performance drops to 58–68% when predicting future cases. They also found that predicting outcomes based solely on judges’ surnames could achieve accuracy of 65%, highlighting potential biases in LJP data.
@@
Zhong et al. (2020a) introduce QAjudge, a reinforcement learning (RL)-based model designed to provide interpretable legal judgements by visualising the prediction process. QAjudge uses a Question Net to iteratively select relevant yes-no questions about case facts, an Answer Net to provide answers and a Predict Net to generate the final judgement. The model aims to minimise the number of questions asked. Evaluated on real-world datasets, QAjudge demonstrated potential in providing reliable and transparent legal judgements.
@@
Xu et al. (2020) propose the Law Article Distillation based Attention Network (LADAN), an end-to-end model addressing the issue of confusing charges in LJP by distinguishing similar law articles. The model uses a novel graph NN to learn differences between confusing law articles and an attention mechanism to extract discriminative features from fact descriptions. Experiments on real-world datasets showed that LADAN improved performance over previous methods in law article prediction, charge prediction and penalty term prediction.
@@
Ma et al. (2021) introduce MSJudge, a MTL framework designed to predict legal judgements by leveraging multi-stage judicial data, including pre-trial claims and court debates. MSJudge consists of components to encode multi-stage context, model interactions among claims, facts and debates and predict judgements. Evaluated on a large civil trial dataset, MSJudge more accurately characterises the interactions among claims, facts and debates for judgement prediction, achieving improvements over state-of-the-art (SOTA) baselines.
@@
Feng et al. (2022) address limitations of SOTA LJP models by proposing an event-based prediction model with constraints to improve performance. The model extracts fine-grained key events from case facts and predicts judgements based on these events rather than the entire fact statement. They manually annotated a legal event dataset and introduced output constraints to guide learning. Their method leverages event information and cross-task consistency constraints.
@@
Tong et al. (2024) introduce GJudge, a graph boosting framework incorporating constraints to address shortcomings of traditional LJP methods. GJudge features a multi-perspective interactive encoder and a Multi-Graph Attention Network (MGAT) consistency expert module. The encoder integrates fact descriptions with label similarity connections, while the expert module differentiates similar labels and preserves task consistency. Experiments on the CAIL datasets show that GJudge outperforms other models, including the SOTA RLJP (Wu et al., 2022), achieving higher F1 scores.
@@
Previous works mainly focused on creating accurate representations of a case’s fact description to enhance judgement prediction performance. However, these methods often overlook the practical judicial process, where human judges compare similar law articles or potential charges before making a decision. To address this gap, Zhang et al. (2023a) propose CL4LJP, a supervised contrastive learning framework to improve LJP by capturing fine-grained differences between similar law articles and charges. The framework includes contrastive learning tasks at the article, charge and label levels, enhancing the model’s ability to model relationships between fact descriptions and labels.
@@
Liu et al. (2023) propose ML-LJP, a multi-law aware LJP method that expands law article prediction into a multi-label classification task incorporating both charge-related and term-related articles. The approach uses label-specific representations and contrastive learning to distinguish similar definitions. A Graph Attention Network (GAT) is employed to learn interactions among multiple law articles for prison term prediction. Experiments showed that ML-LJP outperformed SOTA models, particularly in prison term prediction.
@@ 7. Legal Text Classification
LTC is an important task within the domain of NLP that involves categorising legal documents based on their content, a foundational aspect of building intelligent legal systems. With the exponential growth of legal documents, it has become increasingly challenging for legal professionals to locate relevant rulings in similar cases for argumentation. LTC addresses this challenge by automatically associating legal texts with predefined categories, such as criminal, civil or administrative cases, thereby simplifying legal research and decision-making processes.
@@
In the legal field, this process is often referred to as predictive coding, where ML algorithms are trained through supervised learning to classify documents into specific categories. The broader task of text classification in NLP involves assigning one or multiple categories to a document from a set of predefined options and it can take various forms, including binary classification (predicting whether a document is in a given class or not), multi-class classification (predicting which one of many classes a document belongs to) and multi-label classification (predicting which set of classes a document belongs to). Legal document classification often falls under large multi-label text classification, where the label space can consist of thousands of potential categories, adding complexity to the task (Shaheen et al., 2020).
@@ 7.1. Datasets
Chalkidis et al. (2019b) release EURLEX57K, a dataset containing 57,000 EU legislative documents from the EUR-Lex portal [^4] https://eur-lex.europa.eu, annotated with EuroVoc [^5] https://op.europa.eu/en/web/eu-vocabularies concepts. This dataset facilitates research in LTC, including extreme multi-label text classification, few-shot and zero-shot learning, with documents tagged with an expansive set of descriptors.
Tuggener et al. (2020) introduce LEDGAR, a multi-label corpus of legal provisions from contracts scraped from the the US Securities and Exchange Commission’s website. The dataset includes over 846,000 provisions across 60,540 contracts, with an extensive label set suitable for text classification and legal studies.
@@
Papaloukas et al. (2021) introduce the Greek Legal Code dataset, categorising approximately 47,000 Greek legislative documents into a detailed multi-level classification system. The dataset is structured into volumes, chapters and subjects, each containing diverse legal documents from Greek legislation history, supporting LTC in the Greek legal domain.
@@
Song et al. (2022) introduce POSTURE50K, a legal dataset containing 50,000 the US legal opinions annotated with Legal Procedural Postures ranging from common to rare motions. The dataset includes an innovative split strategy to support supervised and zero-shot learning evaluations, ensuring infrequent categories are adequately represented, enhancing model generalizability and testing accuracy.
@@
Graham et al. (2023) develop a domain-specific dataset for LTC that focuses on deontic modalities in contract sentences. They manually annotated sentences from the CUAD dataset (Hendrycks et al., 2021) for permissions, obligations and prohibitions, providing a resource for modelling and studying these functional categories in legal analysis.
@@
Galassi et al. (2024) extend the Claudette [^6] https://claudette.eui.eu corpus from 25 to 50 ToS per language (English, German, Italian, and Polish). Each sentence is labelled according to nine categories of potential unfairness, with annotations indicating the degree of unfairness. The dataset was carefully compiled based on language availability, structural similarity and version correspondence. Cross-lingual analysis revealed notable differences between languages – particularly in German – including variations in length, structure, missing clauses, and legal terminology that reflect manual drafting adjustments rather than simple automated translations.
7.2. Approaches
DL methods typically requires extensive data to yield effective results, but MTL can help mitigate data scarcity. Elnaggar et al. (2018) leverage transfer learning and MTL to perform tasks such as translation and multi-label classification within legal document corpora. They employ theMultiModel algorithm (Kaiser et al., 2017), a fully convolutional sequence-to-sequence architecture that integrates multiple modality networks. The model maps legal texts into a shared embedding space, enabling task switching and improving generalisation across tasks to make efficient use of limited legal data.
Lee and Lee (2019) examine LTC in Korean by comparing three DL architectures: CNN with ASCII encoding, CNN with Word2Vec embeddings, and RNN with Word2Vec embeddings. Each model assigns case documents to civil, criminal, or administrative categories. Using a dataset of nearly 60,000 past case documents, the study finds that the RNN model with Word2Vec embedding achieves the highest classification accuracy.
Bambroo and Awasthi (2021) introduce an architecture that integrates long attention mechanisms with a distilled BERT model pre-trained on legal domain-specific corpora. Their model employs a combination of local windowed attention and task-motivated global attention to handle inputs up to eight times longer than standard BERT models. The architecture, based on the lightweight DistilBERT transformer (Sanh et al., 2020), and incorporating LongformerSelf-Attention, is optimised for legal document classification, outperforming a fine-tuned BERT model and other transformer-based models in both speed and performance.
Song et al. (2022) present a DL-based system built on top of RoBERTa (Liu et al., 2019) for multi-label legal document classification. They enhance the model with domain-specific pre-training, a label-attention mechanism and MTL to improve classification accuracy, particularly for low-frequency classes. The label-attention mechanism uses label embeddings to bridge the semantic gap between samples and class labels, addressing class imbalance issues.
Wang et al. (2022) introduce a Document-to-Graph Classifier to classify legal documents based on facts and reasons rather than topics. They extract key entities and represented legal documents using four distinct relation graphs capturing different aspects of entity relationships. A graph attention network (Veličković et al., 2018) is used to learn document representations from the combined graph, improving classification by focusing on factual content.
Mamooler et al. (2022) propose an active learning pipeline to fine-tune PLMs for LTC, thereby addressing the challenges of a specialised vocabulary and high annotation costs. Their method involves continual pre-training of RoBERTa on legal texts, knowledge distillation using a pre-trained sentence transformer, and an initial sampling strategy based on clustering unlabelled data. This approach reduces the number of labelling actions required for LTC tasks, reducing the overall cost of the LTC process.
Grabmair et al. (2015) introduce LUIMA, a system designed for conceptual legal document retrieval, focusing on vaccine injury decisions. LUIMA employs a multi-level classification pipeline: rule-based sub-sentence annotations tag legal concepts, such as terms and mentions, while ML classifiers categorise sentences into argumentative roles, such as legal rules or evidence-based findings. These annotations feed into a sentence-level indexing process using Apache Lucene, enabling semantic querying beyond traditional keyword search. A learning-to-rank module refines the retrieved results by leveraging hand-crafted features such as sentence match counts and term similarity scores.
Galassi et al. (2024) explore the binary classification task of detecting potentially unfair clauses in online Terms of Service (ToS) at the sentence level. Sentences are labelled as either unfair, potentially unfair, or fair. Starting from an English-trained model, the authors compare four cross-lingual approaches: (1) training separate models on each language, (2) projecting annotations from English to another language, (3) translating English training documents and using their original annotations and (4) translating query documents into English during prediction. The evaluation shows scenario (4) achieves similar or better results than scenario (1) when translation quality is high. If the translation quality is low, scenario (1) can produce better results. Scenarios (2) and (3) perform slightly below scenario (1) but remain viable alternatives.
Rhetorical role labelling is the task of classifying sentences in legal documents based on their functional roles, such as fact, argument or ruling, to structure and analyse judicial decisions. The Artificial Intelligence for Legal Assistance (AILA) shared task series focuses on advancing legal NLP by introducing datasets and challenges that address core legal text processing tasks. The latest edition, AILA 2021 (Parikh et al., 2022), featured a rhetorical role labelling task that required classifying sentences into seven predefined roles. Building on these challenges, DeepRhole (Bhattacharya et al., 2023) introduces a transformer-based framework that fine-tunes domain-adapted models such as LegalBERT to capture the semantics of legal texts. The system applies different word embedding strategies, including Law2Vec and embeddings from Google News, to model legal language variability. An inter-annotator study examines the subjectivity in rhetorical role assignment, further assessing the model’s performance in different judicial contexts.